OpenStudy (anonymous):

Find the inverse of the function. f(x)=8x+2 A)f-1(x)=log8(x+2) B)f-1(x)=log8(x+8) C)f-1(x)=log8(x-8) D)f-1(x)=log8(x-2)

OpenStudy (aum):

\[ f(x) = 8^x + 2\\ y = 8^x + 2\\ x = 8^y + 2\\ x - 2 = 8^y \text{Take log to the base 8 on both sides:}\\ \log_8(x-2) = y \\ f^{-1}(x) = \log_8(x-2) \]

OpenStudy (anonymous):

thank you

OpenStudy (aum):

yw.